OpenStudy (anonymous):

Solve for x. 7x + 6 = 5x + 30 18 -18 -12 12

OpenStudy (anonymous):

7x+6=5x+30 7x-5x=30-6 2x=24 so x=12 Understand?

OpenStudy (anonymous):

d.12